# Towards Real-World Stickers Use: A New Dataset for Multi-Tag Sticker Recognition

### v2

@@ 6 Conclusion
Stickers can be interpreted differently by users in the real-world scenario, leading to a variety of associated tags. In this paper, we introduce the StickerTAG dataset, comprising 461 tags and 13,571 sticker-tag pairs. Additionally, we present the AttPL framework, specifically designed to tackle the distinct challenges of multi-tag sticker recognition. In detail, we first obtain the attribute-oriented descriptions from stickers based on four attributes. Next, we introduce a local re-attention module to focus on local information, followed by the application of prompt learning to guide the recognition process. Furthermore, we utilize confidence penalty optimization to penalize the confident output distributions. Our AttPL method outperforms existing methods on our StickerTAG dataset, emphasizing the necessity of incorporating stickers’ distinctive characteristics into recognition tasks.

### v3

@@ 6 Conclusion
Stickers can be interpreted differently by users in the real-world scenario, leading to a variety of associated tags. This paper introduces a StickerTAG dataset, comprising 461 tags and 13,571 sticker-tag pairs. Additionally, we present the Att2PL framework, specifically designed to tackle the distinct challenges of multi-tag sticker recognition. We first obtain the attribute-oriented descriptions from stickers based on four attributes. Next, we introduce a local re-attention module to focus on local information, followed by the application of prompt learning to guide the recognition process. Furthermore, we utilize confidence penalty optimization to penalize the confident output distributions. Our Att2PL method outperforms existing methods on our StickerTAG and SER30K dataset, emphasizing the necessity of incorporating stickers’ distinctive characteristics into recognition tasks.